OpenStudy (anonymous):

Plot .(my three key doesnt work..) -x+three y=-6

OpenStudy (phi):

-x+3y= -6 this is the equation of a line one way to plot is find two (x,y) pairs, plot them and draw a line through the 2 points if x=0, what is y? if y is 0, what is x? answer those two questions to get 2 (x,y) pairs

OpenStudy (anonymous):

ok so -6= -x+three y

OpenStudy (phi):

that is the exact same equation. now replace x with 0. what do you get?

OpenStudy (anonymous):

-6=-0+three y

OpenStudy (anonymous):

do i subtract three y both sides

OpenStudy (phi):

-6= -0+3y of course -0 is the same as 0 (0 is 0) and 0 added to anything does not change anything: -6= 3y to find y, divide both sides by 3

OpenStudy (anonymous):

x=three

OpenStudy (anonymous):

y=three

OpenStudy (phi):

you have to be more careful. first, there is no x. it's y 2nd, -6/3 = -2 not +2 or +3 3y= -6 3y/3 = -6/3 y= -2

OpenStudy (phi):

that means (0,-2) is on the line now replace y with 0 in -x+3y= -6

OpenStudy (anonymous):

ok i do see the mistake x but its y=

OpenStudy (anonymous):

-x+three(-2)=-6

OpenStudy (anonymous):

i think i put th

OpenStudy (phi):

no, we are done with the first part of the problem. now replace y with 0 in -x+3y= -6

OpenStudy (anonymous):

the wrong number for x ok. -x+three(0)=-6

OpenStudy (phi):

yes -x +3*0= -6 -x+0 =-6 -x= -6 if you multiply (or divide!) by -1 on both sides: -1* -x= -1*-6 what do you get?

OpenStudy (anonymous):

(0,6)

OpenStudy (anonymous):

ok or x=6

OpenStudy (phi):

no, be more careful -1* -x= -1*-6 what do you get? you get x= 6 that means the pair is (6,0) you now have 2 points to plot (0,-2) and (6,0)

OpenStudy (anonymous):

i have trouble with that swapping numbers

OpenStudy (phi):

if you know that, write down (x,y) on your paper and double check with this last one, we started with y=0 so write down (x,0) so you don't forget then work out what x is , and fill it in: (6,0)

OpenStudy (anonymous):

ok so th value is -three y and 6x

OpenStudy (anonymous):

ok that i will do to help me

OpenStudy (phi):

you now have 2 points to plot (0,-2) and (6,0)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok did i get my value correct

OpenStudy (phi):